Write first five multiples of: 
5 ???????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The required multiples are:
5 ×1 =5
5 × 2 = 10
5 × 3 = 15
5 × 4 = 20
5 × 5 = 25
Hence, the first five multiples of 5 are 5, 10, 15, 20 and 25

Thanks